OpenStudy (anonymous):

is the derivative function of: f(x)= sqrt[4]{x} cosx help please...

OpenStudy (anonymous):

Please write it claerly. Is the Sqrt is common?

OpenStudy (anonymous):

question is not clear!

OpenStudy (anonymous):

\[f(x)= \sqrt[4]{x} cosx\]

OpenStudy (experimentx):

use chain rule \[ f'(x) = \frac{d (x \cos (x))^{1/4}}{d (x \cos (x))} \times { d (x\cos (x)) \over dx}\]

OpenStudy (unklerhaukus):

apply te product rule \[\left(gh\right)^\prime=gh^\prime+g^\prime h\] \[f(x)=\sqrt[4]x\cos x=x^{1/4}\cos x\] \[=-x^{1/4}\sin x+\frac 14x^{-3/4}\cos x\] \[=-\sqrt[4]x\sin x+\frac {\cos x}{4x^{3/4}}\]